OpenStudy (anonymous):

help me to solve this question... Find the total area between the curve y=1-x^2, and the x-axis over the interval (0,2)...

OpenStudy (anonymous):

you have to integrate the given curve's equation in the interval (0,2) \[\int\limits_{0}^{2}1-x^2 = (2-0) + (-8/3 + 0) = 2 - 8/3 = -2/3\] Since are cannot be negative, area = 2/3 sq.units

OpenStudy (anonymous):

sorry I'm little confused while typing equations on OpenStudy..so I've been a little late in answering

OpenStudy (anonymous):

oo its ok..the working just like that AbhimanyuPudi ?

OpenStudy (anonymous):

may be u have to put another step in between..showing the integration.. \[\int\limits_{0}^{2}1-x^2 = x(from 0 \to 2) - x^3/3 (from 0 \to 2)\]

OpenStudy (anonymous):

the negative sign in the answer indicates that the area is formed below the x-axis

OpenStudy (raden):

for this case, first u have to figure out of the function y=1-x^2  so, the total area = A1 + A2